OpenStudy (kittykat):

Find the vector and parametric equations of the following lines. (c) The line passing through P(3, −1, 4) and Q(3, −1, 5). (e) The line passing through P(1, 0, −3) and parallel to the line with parametric equations x = −1 + 2t, y = 2 − t, z = 3 + 3t.

Parth (parthkohli):

Hmm, alright. Where can we start with this?

Parth (parthkohli):

Some starters... The direction vector is the vector that joins any two points on a line. A vector lying on the line is any multiple of the direction vector added to a point lying on the line.

OpenStudy (kittykat):

@3mar can you help me ?

OpenStudy (3mar):

It is not from honesty to answer a question I do not know. I wish I could have been more help. Sorry I am not so good at it!!

OpenStudy (irishboy123):

vector equation can be written as \(\vec r = \vec r_o + t \vec d\) \(\vec r_o\) is any point on the line, \(\vec d\) is the line's direction vector, t is parameter

OpenStudy (irishboy123):

so \(\vec r_o\) can be either of (3, −1, 4) or (3, −1, 5) [or any other point on the line but you're given those ones and direction can be \(\vec d = 3, −1, 4) - (3, −1, 5)\), to start with at least, you can scale it down as long as you keep its direction

OpenStudy (dumbcow):

To finish this question, let me continue what has been previously said: \[d = (3,-1,5) - (3,-1,4) = (0,0,1)\] vector equation is: \[r(t) = (3,-1,4) + (0,0,1)t\] parametric equation is: \[x(t) = 3, y(t) = -1, z(t) = 4+t\] Part e) Parallel lines have the same directional vector ---> (2,-1,3) vector equation: \[r(t) = (1,0,-3) + (2,-1,3)t\] parametric equation: \[x(t) = 1+2t, y(t) = -t, z(t) = -3+3t\]